OpenStudy (anonymous):

What is the 14th term for the geometric sequence whose first term is 12 and whose ratio is 1/2

OpenStudy (anonymous):

geometric: ar^n we have r = 1/2 also: 12 = a(1/2)^0 which is 12 = a plug, using n = 14

OpenStudy (anonymous):

3/4096 is answer or no?

OpenStudy (anonymous):

you got it.